what is the foci of the hyperbola with the equation y²/144-x²/100=1
Answer by Nate(3500)

find the foci of the hyperbola with equation y²/144-x²/100=1
vertical transverse axis
center(0,0)
a=12
b=10
c^2=a^2+b^2
c^2=144+100
c^2=244
c=(244)^(1/2)=4(61)^(1/2)
Foci: (0,4(61)^(1/2)),(0,-4(61)^(1/2))